You randomly select one card from a deck of 52 cards. Find the probabilty of selecting an ace or a 8?
Found 2 solutions by Mathtut, stanbon:
Answer by Mathtut(3670)

there are 4 aces therefore the probability is 4/52=1/13
there are 4 8's therefore the probability would be the same 1/13
:
if aces and 8's together it would be 8/52=2/13

Answer by stanbon(75887)

You randomly select one card from a deck of 52 cards. Find the probabilty of selecting an ace or a 8?
--------------------------
# of ways to succeed: 8
# of possible outcomes: 52
P(ace or 8) = 8/52 = 2/13 = 0.153846...
Cheers,
Stan H.